# Manifold Learning by Mixture Models of VAEs for Inverse Problems

### v2

@@ 1 Introduction
The treatment of high-dimensional data is often computationally costly and numerically unstable. Therefore, in many applications, it is important to find a low-dimensional representation of high-dimensional datasets. Classical methods, like the principal component analysis (PCA, [80]), assume that the data is contained in a low-dimensional subspace. However, for complex datasets this assumption appears to be too restrictive, particularly when working with image datasets. Therefore, recent methods rely on the so-called manifold hypothesis ([16]), stating that even complex and high-dimensional datasets are contained in a low-dimensional manifold. Based on this hypothesis, in recent years, many successful approaches have been based on generative models, able to represent high dimensional data in $\mathbb{R}^{n}$ by a generator $D\colon\mathbb{R}^{d}\to\mathbb{R}^{n}$ with $d\ll n$: these include generative adversarial networks (GANs, [42]), variational autoencoders (VAEs, [61]), injective flows ([67]) and score-based diffusion models ([91, 52]). For a survey on older approaches to manifold learning, the reader is referred to [72, 57] and to the references therein.
@@
Under the assumption that $D$ is injective, the set of generated points $\{D(z):z\in\mathbb{R}^{d}\}$ forms a manifold that approximates the training set. However, this requires that the data manifold admits a global parameterization. In particular, it must not be disconnected or contain holes. In order to model disconnected manifolds, [36, 58, 64, 81] propose to model the latent space of a VAE by a Gaussian mixture model. This enables the authors to capture multimodal probability distributions. However, this approach struggles with modelling manifolds with holes since either the injectivity of the generator is violated or it is impossible to model overlapping charts. Similarly, the authors of [32, 75, 83] propose latent distributions defined on Riemannian manifolds for representing general topologies. In [74], the manifold is embedded into a higher-dimensional space, in the spirit of Whitney embedding theorem. However, these approaches have the drawback that the topology of the manifold has to be known a priori, which is usually not the case in practice.
@@
Here, we focus on the representation of the data manifold by several charts. A chart provides a parameterization of an open subset from the manifold by defining a mapping from the manifold into a Euclidean space. Then, the manifold is represented by the collection of all of these charts, which is called atlas. For finding these charts, the authors of [30, 40, 82, 88] propose the use of clustering algorithms. By default, these methods do not provide an explicit formulation of the resulting charts. As a remedy, the authors of [21, 82] use linear or kernelized embeddings. The authors of [88] propose to learn for each chart again a generative model. However, these approaches often require a large number of charts and are limited to relatively low data dimensions. The idea of representing the charts by generative models is further elaborated in [65, 66, 86]. Here, the authors proposes to train at the same time several (non-variational) autoencoders and a classification network that decides for each point to which chart it belongs. In contrast to the clustering-based algorithms, the computational effort scales well for large data dimensions. On the other hand, the numerical examples in the corresponding papers show that the approach already has difficulties to approximate small toy examples like a torus.
@@
In this paper, we propose to approximate the data manifold by a mixture model of VAEs. Using Bayes theorem and the ELBO approximation of the likelihood term we derive a loss function for maximum likelihood estimation of the model weights. Mixture models of generative models for modelling disconnected datasets were already considered in [13, 53, 69, 92, 94]. However, they are trained in a different way and to the best of our knowledge none of those is used for manifold learning.
@@
where $\mathcal{G}\colon\mathbb{R}^{n}\to\mathbb{R}^{m}$ is an ill-posed or ill-conditioned, possibly nonlinear, forward operator and $\eta$ represents additive noise. Reconstructing the input $x$ directly from the observation $y$ is usually not possible due to the ill-posed operator and the high dimension of the problem. As a remedy, the incorporation of prior knowledge is required. This is usually achieved by using regularization theory, namely, by minimizing the sum of a data fidelity term $F(x)$ and a regularizer $R(x)$, where $F$ describes the fit of $x$ to $y$ and $R$ incorporates the prior knowledge. With the success of deep learning, data-driven regularizers became popular ([6, 9, 43, 51, 71]).
@@
where $F(x)=\frac{1}{2}\|\mathcal{G}(x)-y\|^{2}$ is a data-fidelity term. This corresponds to the regularizer $R(x)$ which is zero for $x\in\mathcal{M}$ and infinity otherwise. When the manifold admits a global parameterization given by one single generator $D$, the authors of [4, 24, 34, 41] propose to reformulate the problem as $\hat{x}=\hat{z}$, where $\hat{z}=\operatorname*{arg\,min}_{z}F(D(x))$. Since this is an unconstrained problem, it can be solved by gradient based methods. However, since we consider manifolds represented by several charts, this reformulation cannot be applied. As a remedy, we propose to use a Riemannian gradient descent scheme. In particular, we derive the Riemannian gradient using the decoders and encoders of our manifold and propose two suitable retractions for applying a descent step into the gradient direction.
To emphasize the advantage of using multiple generators, we demonstrate the performance of our method on numerical examples. We first consider some two- and three-dimensional toy examples. Finally, we apply our method to deblurring and to electrical impedance tomography (EIT), a nonlinear inverse problem consisting in the reconstruction of the leading coefficient of a second order elliptic PDE from the knowledge of the boundary values of its solutions ([28]). The code of the numerical examples is available online [^1] https://github.com/johertrich/Manifold_Mixture_VAEs.
@@ 2.1 Variational Autoencoders for Manifold Learning
In this paper, we assume that we are given data points $x_{1},\dots,x_{N}\in\mathbb{R}^{n}$ for a large dimension $n$. In order to reduce the computational effort and to regularize inverse problems, we assume that these data-points are located in a lower-dimensional manifold. We aim to learn the underlying manifold from the data points $x_{1},\dots,x_{N}$ with a VAE ([61, 62]).
@@
The log-density $\log(p_{\tilde{X}}(x))$ induced by the model is called the evidence. However, for VAEs the computation of the evidence is intractable. Therefore, the authors of [61] suggest to approximate it by the evidence lower bound given by
@@
It is a known issue of VAEs that the inferred probability distribution is often more blurry than the ground truth distribution of the data. A detailed discussion of this issue can be found in [62, Section 2.8.2]. As a remedy, the authors suggest to choose a more flexible model. One possibility is to combine VAEs with normalizing flows, as proposed in [84] or in [31]. Following these approaches, we increase the flexibility of the model by using a latent space learned by a normalizing flow. The idea is based on the observation that transforming probability distributions in low-dimensional spaces is much cheaper than in high-dimensional spaces. Consequently, modelling the low-dimensional latent space can be more effective than learning probability transformations in the high-dimensional data space. Here, we employ the specific loss function from [45, 46] for training the arising model. More precisely, we choose the latent distribution
@@
In the literature, there exist several invertible neural network architectures based on coupling blocks ([33, 60]), residual networks ([15, 25, 50]), ODE representations ([26, 44, 78]) and autoregressive flows ([54]). In our numerics, we use the coupling-based architecture from [8].
@@
In order to obtain a lower-dimensional representation of the data points, some papers propose to approximate the data-manifold by $\mathcal{M}\coloneqq\{D(z):z\in\mathbb{R}^{d}\}$, see e.g. [4, 24, 34, 41]. However, this is only possible if the data-manifold admits a global parameterization, i.e., it can be approximated by one generating function. This assumption is often violated in practice. As a toy example, consider the one-dimensional manifold embedded in $\mathbb{R}^{2}$ that consists of two circles, see Figure 1(a). This manifold is disconnected and contains “holes”. Consequently, the topologies of the manifold and of the latent space $\mathbb{R}$ do not coincide, so that the manifold cannot be approximated by a VAE. Indeed, this can be verified numerically. When we learn a VAE for approximating samples from this manifold, we observe that the two (generated) circles are not closed and that both components are connected, see Figure 1(b). As a remedy, in the next section, we propose the use of multiple generators to resolve this problem, see Figure 1(c).For this purpose, we need the notion of charts and atlases.
@@ 3.1 Training of Mixtures of VAEs
Let $x_{1},\dots,x_{N}$ be the noisy training samples. In order to train mixtures of VAEs, we again minimize an approximation of an upper bound to the negative log likelihood function $-\sum_{i=1}^{N}\log(p_{\tilde{X}}(x_{i}))$. To this end, we employ the law of total probability and Jensen’s inequality to obtain
$\log(p_{\tilde{X}}(x_{i}))\geq\sum_{k=1}^{K}\beta_{ik}\log(p_{\tilde{X}_{k}}(x_{i})),$
where $\beta_{ik}\coloneqq P(J=k|\tilde{X}=x_{i})$ is the probability that the sample $x_{i}$ was generated by the $k$-th random variable $\tilde{X}_{k}$. Using the definition of conditional probabilities, we observe that
@@
$\log(p_{\tilde{X}}(x_{i}))\approx\ell(x_{i}|\Theta)=\sum_{k=1}^{K}\tfrac{\alpha_{k}\exp(\mathrm{ELBO}(x_{i}|\theta_{k}))}{\sum_{j=1}^{K}\alpha_{j}\exp(\mathrm{ELBO}(x_{i}|\theta_{j}))}\mathrm{ELBO}(x_{i}|\theta_{k}).$ (6)
@@
in order to train the parameters $\Theta=(\theta_{k})_{k=1}^{K}$ of the mixture of VAEs. Finally, this loss function is then optimized with a stochastic gradient based optimizer like Adam ([59]).
Remark 1 (Lipschitz regularization).
@@
using a certain number of epochs of the Adam optimizer ([59]).
@@
Remark 2 (Number of charts $K$).
@@ 3.2 Architectures
In this paper, we build the invertible neural networks $T_{l}$ and the normalizing flows $\mathcal{T}_{k}$ out of coupling blocks as proposed in [8] based on the real NVP architecture ([33]). To this end, we split the input $z\in\mathbb{R}^{d_{l}}$ into two parts $z=(z_{1},z_{2})\in\mathbb{R}^{d_{l}^{1}}\times\mathbb{R}^{d_{l}^{2}}$ with $d_{l}=d_{l}^{1}+d_{l}^{2}$ and define $T_{l}(z)=(x_{1},x_{2})$ with
@@
Remark 3 (Projection onto learned charts).
@@ 4 Optimization on Learned Manifolds
As outlined in the previous section, the inverse charts $\varphi_{k}^{-1}$ of the manifold $\mathcal{M}$ are modeled by $\mathcal{D}_{k}\coloneqq D_{k}\circ\mathcal{T}_{k}$. The chart $\varphi_{k}$ itself is given by the mapping $\mathcal{E}_{k}\coloneqq\mathcal{T}_{k}^{-1}\circ E_{k}$ restricted to the manifold. For the special case of a VAE with a single generator $\mathcal{D}$, the authors of [4, 24, 34] propose to solve (12) in the latent space. More precisely, starting with a latent initialization $z_{0}\in\mathbb{R}^{d}$ they propose to solve
@@
Therefore, we aim to use a gradient formulation which is independent of the parameterization of the manifold. Here, we use the concept of the Riemannian gradient with respect to the Riemannian metric, which is inherited from the Euclidean space in which the manifold $\mathcal{M}$ is embedded. To this end, we first revisit some basic facts about Riemannian gradients on embedded manifolds which can be found, e.g., in [1]. Afterwards, we consider suitable retractions in order to perform a descent step in the direction of the negative Riemannian gradient. Finally, we use these notions in order to derive a gradient descent procedure on a manifold given by mixtures of VAEs.

### v3

@@ Abstract
Keywords: manifold learning, mixture models, variational autoencoders, Riemannian optimization, inverse problems
@@ 1 Introduction
The treatment of high-dimensional data is often computationally costly and numerically unstable. Therefore, in many applications, it is important to find a low-dimensional representation of high-dimensional data sets. Classical methods, like the principal component analysis (PCA, Pearson, 1901), assume that the data is contained in a low-dimensional subspace. However, for complex data sets this assumption appears to be too restrictive, particularly when working with image data sets. Therefore, recent methods rely on the so-called manifold hypothesis (Bengio et al., 2013), stating that even complex and high-dimensional data sets are contained in a low-dimensional manifold. Based on this hypothesis, in recent years, many successful approaches have been based on generative models, able to represent high dimensional data in $\mathbb{R}^{n}$ by a generator $D\colon\mathbb{R}^{d}\to\mathbb{R}^{n}$ with $d\ll n$: these include generative adversarial networks (GANs, Goodfellow et al., 2014), variational autoencoders (VAEs, Kingma and Welling, 2014), injective flows (Kothari et al., 2021) and score-based diffusion models (Song and Ermon, 2019; Ho et al., 2020). For a survey on older approaches to manifold learning, the reader is referred to Ma and Fu (2011); Izenman (2012) and to the references therein.
@@
Under the assumption that $D$ is injective, the set of generated points $\{D(z):z\in\mathbb{R}^{d}\}$ forms a manifold that approximates the training set. However, this requires that the data manifold admits a global parameterization. In particular, it must not be disconnected or contain holes. In order to model disconnected manifolds, Falck et al. (2021); Jiang et al. (2017); Kivva et al. (2022); Pineau and Lelarge (2018) propose to model the latent space of a VAE by a Gaussian mixture model. This enables the authors to capture multimodal probability distributions. However, this approach struggles with modelling manifolds with holes since either the injectivity of the generator is violated or it is impossible to model overlapping charts. Similarly, Davidson et al. (2018); Mathieu et al. (2019); Rey et al. (2020) propose latent distributions defined on Riemannian manifolds for representing general topologies. Massa et al. (2022) embed the manifold into a higher-dimensional space, in the spirit of Whitney embedding theorem. However, these approaches have the drawback that the topology of the manifold has to be known a priori, which is usually not the case in practice.
@@
Here, we focus on the representation of the data manifold by several charts. A chart provides a parameterization of an open subset from the manifold by defining a mapping from the manifold into a Euclidean space. Then, the manifold is represented by the collection of all of these charts, which is called atlas. For finding these charts, Cohn et al. (2021); Floryan and Graham (2022); Pitelis et al. (2013); Sidheekh et al. (2022) propose the use of clustering algorithms. By default, these methods do not provide an explicit formulation of the resulting charts. As a remedy, Brand (2002); Pitelis et al. (2013) use linear or kernelized embeddings. Sidheekh et al. (2022) propose to learn for each chart again a generative model. However, these approaches often require a large number of charts and are limited to relatively low data dimensions. The idea of representing the charts by generative models is further elaborated by Korman (2018, 2021); Schonsheck et al. (2019). Here, the authors proposes to train at the same time several (non-variational) autoencoders and a classification network that decides for each point to which chart it belongs. In contrast to the clustering-based algorithms, the computational effort scales well for large data dimensions. On the other hand, the numerical examples in the corresponding papers show that the approach already has difficulties to approximate small toy examples like a torus.
@@
In this paper, we propose to approximate the data manifold by a mixture model of VAEs. Using Bayes theorem and the ELBO approximation of the likelihood term we derive a loss function for maximum likelihood estimation of the model weights. Mixture models of generative models for modelling disconnected data sets were already considered by Banijamali et al. (2017); Hoang et al. (2018); Locatello et al. (2018); Stolberg-Larsen and Sommer (2022); Ye and Bors (2022). However, they are trained in a different way and to the best of our knowledge none of those is used for manifold learning.
@@
where $\mathcal{G}\colon\mathbb{R}^{n}\to\mathbb{R}^{m}$ is an ill-posed or ill-conditioned, possibly nonlinear forward operator and $\eta$ represents additive noise. Reconstructing the input $x$ directly from the observation $y$ is usually not possible due to the ill-posed operator and the high dimension of the problem. As a remedy, the incorporation of prior knowledge is required. This is usually achieved by using regularization theory, namely, by minimizing the sum of a data fidelity term $F(x)$ and a regularizer $R(x)$, where $F$ describes the fit of $x$ to $y$ and $R$ incorporates the prior knowledge. With the success of deep learning, data-driven regularizers became popular (Altekrüger et al., 2023; Arridge et al., 2019; Goujon et al., 2023; Hertrich et al., 2022; Lunz et al., 2018).
@@
where $F(x)=\frac{1}{2}\|\mathcal{G}(x)-y\|^{2}$ is a data-fidelity term. This corresponds to the regularizer $R(x)$ which is zero for $x\in\mathcal{M}$ and infinity otherwise. When the manifold admits a global parameterization given by one single generator $D$, Alberti et al. (to appear); Chen et al. (2020); Duff et al. (2021); González et al. (2022) propose to reformulate the problem as $\hat{x}=\hat{z}$, where $\hat{z}=\operatorname*{arg\,min}_{z}F(D(x))$. Since this is an unconstrained problem, it can be solved by gradient based methods. However, since we consider manifolds represented by several charts, this reformulation cannot be applied. As a remedy, we propose to use a Riemannian gradient descent scheme. In particular, we derive the Riemannian gradient using the decoders and encoders of our manifold and propose two suitable retractions for applying a descent step into the gradient direction.
To emphasize the advantage of using multiple generators, we demonstrate the performance of our method on numerical examples. We first consider some two- and three-dimensional toy examples. Finally, we apply our method to deblurring and to electrical impedance tomography (EIT), a nonlinear inverse problem consisting in the reconstruction of the leading coefficient of a second order elliptic PDE from the knowledge of the boundary values of its solutions (Cheney et al., 1999). The code of the numerical examples is available online. [^1] The code is available at https://github.com/johertrich/Manifold_Mixture_VAEs
@@ 2.1 Variational Autoencoders for Manifold Learning
In this paper, we assume that we are given data points $x_{1},\dots,x_{N}\in\mathbb{R}^{n}$ for a large dimension $n$. In order to reduce the computational effort and to regularize inverse problems, we assume that these data-points are located in a lower-dimensional manifold. We aim to learn the underlying manifold from the data points $x_{1},\dots,x_{N}$ with a VAE (Kingma and Welling, 2014, 2019).
@@
The log-density $\log(p_{\tilde{X}}(x))$ induced by the model is called the evidence. However, for VAEs the computation of the evidence is intractable. Therefore, Kingma and Welling (2014) suggest to approximate it by the evidence lower bound given by
@@
It is a known issue of VAEs that the inferred probability distribution is often more blurry than the ground truth distribution of the data. A detailed discussion of this issue can be found in Section 2.8.2 of the survey paper by Kingma and Welling (2019). As a remedy, the authors suggest to choose a more flexible model. One possibility is to combine VAEs with normalizing flows, as proposed by Rezende and Mohamed (2015) or Dai and Wipf (2019). Following these approaches, we increase the flexibility of the model by using a latent space learned by a normalizing flow. The idea is based on the observation that transforming probability distributions in low-dimensional spaces is much cheaper than in high-dimensional spaces. Consequently, modelling the low-dimensional latent space can be more effective than learning probability transformations in the high-dimensional data space. Here, we employ the specific loss function proposed by Hagemann et al. (2022, 2023) for training the arising model. More precisely, we choose the latent distribution
@@
In the literature, there exist several invertible neural network architectures based on coupling blocks (Dinh et al., 2016; Kingma and Dhariwal, 2018), residual networks (Behrmann et al., 2019; Chen et al., 2019; Hertrich, 2023), ODE representations (Chen et al., 2018; Grathwohl et al., 2018; Onken et al., 2021) and autoregressive flows (Huang et al., 2018). In our numerics, we use the coupling-based architecture proposed by Ardizzone et al. (2019).
@@
In order to obtain a lower-dimensional representation of the data points, some papers propose to approximate the data-manifold by $\mathcal{M}\coloneqq\{D(z):z\in\mathbb{R}^{d}\}$ (see, e.g., Alberti et al., to appear; Chen et al., 2020; Duff et al., 2021; González et al., 2022). However, this is only possible if the data-manifold admits a global parameterization, i.e., it can be approximated by one generating function. This assumption is often violated in practice. As a toy example, consider the one-dimensional manifold embedded in $\mathbb{R}^{2}$ that consists of two circles, see Figure 1(a). This manifold is disconnected and contains “holes”. Consequently, the topologies of the manifold and of the latent space $\mathbb{R}$ do not coincide, so that the manifold cannot be approximated by a VAE. Indeed, this can be verified numerically. When we learn a VAE for approximating samples from this manifold, we observe that the two (generated) circles are not closed and that both components are connected, see Figure 1(b). As a remedy, in the next section, we propose the use of multiple generators to resolve this problem, see Figure 1(c). For this purpose, we need the notion of charts and atlases.
@@ 3.1 Training of Mixtures of VAEs
Let $x_{1},\dots,x_{N}$ be the noisy training samples. We initialize the weights $\alpha$ by $\alpha_{k}=\frac{1}{K}$ for all $k$. They will be estimated later in the training algorithm (see Algorithm 1). In order to train mixtures of VAEs, we again minimize an approximation of an upper bound to the negative log likelihood function $-\sum_{i=1}^{N}\log(p_{\tilde{X}}(x_{i}))$. To this end, we employ the law of total probability and Jensen’s inequality to obtain
$\log(p_{\tilde{X}}(x_{i}))$ $=\log\Big{(}\sum_{k=1}^{K}\beta_{ik}p_{\tilde{X}}(x_{i})\big{)}\geq\log\Big{(}\sum_{k=1}^{K}\beta_{ik}\alpha_{k}p_{\tilde{X}_{k}}(x_{i})\Big{)}$ (4)
$\geq\sum_{k=1}^{K}\beta_{ik}\big{(}\log(p_{\tilde{X}_{k}}(x_{i}))+\log(\alpha_{k})\big{)}=\sum_{k=1}^{K}\beta_{ik}\log(p_{\tilde{X}_{k}}(x_{i}))-\log(K)$ (5)
where $\beta_{ik}\coloneqq P(J=k|\tilde{X}=x_{i})$ is the probability that the sample $x_{i}$ was generated by the $k$-th random variable $\tilde{X}_{k}$ and we used that $\alpha_{k}=\frac{1}{K}$. Using the definition of conditional probabilities, we observe that
@@
$\log(p_{\tilde{X}}(x_{i}))\approx\ell(x_{i}|\Theta)=\sum_{k=1}^{K}\tfrac{\alpha_{k}\exp(\mathrm{ELBO}(x_{i}|\theta_{k}))}{\sum_{j=1}^{K}\alpha_{j}\exp(\mathrm{ELBO}(x_{i}|\theta_{j}))}\mathrm{ELBO}(x_{i}|\theta_{k})-\log(K).$ (8)
@@
in order to train the parameters $\Theta=(\theta_{k})_{k=1}^{K}$ of the mixture of VAEs. Finally, this loss function is then optimized with a stochastic gradient based optimizer like Adam (Kingma and Ba, 2015).
Remark 1 (Lipschitz regularization)
@@
using a certain number of epochs of the Adam optimizer (Kingma and Ba, 2015).
@@
Remark 2 (Number of charts $K$)
@@ 3.2 Architectures
In this paper, we build the invertible neural networks $T_{l}$ and the normalizing flows $\mathcal{T}_{k}$ out of coupling blocks as proposed by Ardizzone et al. (2019) based on the real NVP architecture (Dinh et al., 2016). To this end, we split the input $z\in\mathbb{R}^{d_{l}}$ into two parts $z=(z_{1},z_{2})\in\mathbb{R}^{d_{l}^{1}}\times\mathbb{R}^{d_{l}^{2}}$ with $d_{l}=d_{l}^{1}+d_{l}^{2}$ and define $T_{l}(z)=(x_{1},x_{2})$ with
@@
Remark 3 (Projection onto learned charts)
@@ 4 Optimization on Learned Manifolds
As outlined in the previous section, the inverse charts $\varphi_{k}^{-1}$ of the manifold $\mathcal{M}$ are modeled by $\mathcal{D}_{k}\coloneqq D_{k}\circ\mathcal{T}_{k}$. The chart $\varphi_{k}$ itself is given by the mapping $\mathcal{E}_{k}\coloneqq\mathcal{T}_{k}^{-1}\circ E_{k}$ restricted to the manifold. For the special case of a VAE with a single generator $\mathcal{D}$, Alberti et al. (to appear); Chen et al. (2020); Duff et al. (2021) propose to solve (14) in the latent space. More precisely, starting with a latent initialization $z_{0}\in\mathbb{R}^{d}$ they propose to solve
@@
Therefore, we aim to use a gradient formulation which is independent of the parameterization of the manifold. Here, we use the concept of the Riemannian gradient with respect to the Riemannian metric, which is inherited from the Euclidean space in which the manifold $\mathcal{M}$ is embedded. To this end, we first revisit some basic facts about Riemannian gradients on embedded manifolds which can be found, e.g., in the book of Absil et al. (2009). Afterwards, we consider suitable retractions in order to perform a descent step in the direction of the negative Riemannian gradient. Finally, we use these notions in order to derive a gradient descent procedure on a manifold given by mixtures of VAEs.